UT 20140120 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards, DThompson
Telescope Operator: Sallenson
LBTO Support: TSargent, DMiller (for AO-preset test)

Plan for the night

- Test AO-preset
- Image movement due to flexure
- Acq. Telluric for MOS-acq.
- MOS-alignment (training Dave)
- Simulating scientific applications
- Test blind offset


GEIRS stable

Software stable

Deep imaging

One MOS error (MOS-SW, easy fix)

Lots of MOS-acquisition

Fiddling with Telluric acquisition


Check fieldalignment: dx,dy=5600,-1700


All times are UT, files are luci2.20140119.iiii

00:39 Darks for CVeillet, 10*(5*20sec), #9-18, LIR integrated, 10*(10*10sec) LIR integrated #19-28

00:43 Sunset, sky mostly covered, low wind, Hum=20&, T=1deg

01:15 Opened

01:25 Field alignment, #29-38

01:45 Steve checks pointing and collimation

01:52 BS9107, LS-acq-2stars, sky/src #39/40, slit #41, aligned, #42 verification,
Pos star1 requested: x,y=1010/1023, offset=8.7"/-43.2/-35.1deg, star found: x,y=1041/1032,
offset, dx,dy=-31,9, use #42 as input for another alignment (incl. rotation),
does not work??, should do everything in pix-space, but apparently uses info (PA) from header
Redo with just translation, ?? gives also wrong values

02:55 HH91, pretty clear now, #44 FeII, seeing on image 0.7", H2 #45-80 seeing 0.5-0.7", Br_gam #81-116
Duty Cycle 70% (for a script with 73 obsitems, 9*10sec integrated per position and large offsets for sky images)

05:34 Ima movement M67, PA=0,90,180,270, # 117-120 , error FW2 while pointing

05:36:20 *Error message while moving pointing mask in FPU: ...Severe problem detected during motion of motor card number 1, motor number 7...*
Mask was properly put in though

05:55 M67_MOS, Dave's training MOS-acquisiton, files #121-124
Bug in MOS-acq., if one declicks a star when aligning (since its bad) and new calculates new solution,
the next star below has a high error

Tests AO-preset

06:30 Steve checks poining and collimate at a region close to the AO-target, Doug is preparing secondary/WFS

06:43 Alexander did AO-preset via script, not successfull, R/I-mag of AO-reference star must be provided in script

07:00 Dougs does an AO-active preset manually, image of 14th mag star before AO-correction with z+HeI crossed filters #125,
but loop can not be closed, requires some closed-dome tests tomorrow (fiber-problem), postponed for tomorrow night

07:25 Ima movement M67, PA=0,90,180,270, # 126-129

07:50 MOS-acq Sl01, #130-133

08:09 Blind offset, #134-138, OK, worked, bug in scripting tool (right mirror tag)

08:38 MOS-acq. SL03, #, had to stop/restart telescope scheduler, OK, #139-161

09:20 Telluric-acq. with MOS-mask assuming telluric is in same field, #162-163 OK Telluric-acq. on BS9126, #164, no, doesn't work, bug in RTD-too,

09:40 Ima movement M67, PA=0,90,180,270, #165-169

10:21 Jitter-test, RG1312, Ks, #170-189, seeing 0.6-0.7"

11:02 MOS-acq, GDNPHB21, GS is double star, can't do it, no other suitable

11:17 MOS-acq, SDSS1110, MOS-panic at 11:26:39, recovered by Peter
no mechanical reason for fail, minor MOS-SW issue

12:20 Ingo has provided a new RTD-version (better Telluric alignment)

12:24 Reacquire SDSSJ1110, error FW2, recovered, cont'd, didn't work, header keyword have changed tonight
and RTD requires some which are not any longer existent, Alexander fixes it
sky #193, src #194 ,MOSview #195, verification #196, spectra #197-200

13:34 Telluric, Ingos new tool didn't work since some more header keywords are missing,
will be an easy fix tomorrow, did it manually, #201-222 (last 12 frames are Telluric spectra)

14:00 Closing


Mask SDSSJ1110, G150@2.245mue, clear+K, CalUnit in, all images LIR, all normal
Lamp off, 5*2.75s, #223-227
Halo2 on, 5*10sec, #228-232
Lamps off, 5*10sec, #233-237
Argon on, 1*10sec, #238
Neon on, 1*10sec, #239
Xenon on, 1*10sec, #240/241
Lamps off, 5*10sec, #242-246
Spectro sieve, 10sec, G150@2.245mue, clear+K, #247
Spectro sieve, 10sec, G210@2.225mue, clear+K, #248

-- JochenHeidt - 20 Jan 2014
Topic revision: r9 - 21 Jan 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback